you speed up from rest in the forward horizontal direction with a constant magnitude of acceleration equal to 3.4m/s^2 and then some tome later slow down with constant magnitude od acceleration equal to 5m/s^2. Describe and calculate the following for a 0.3 m long dangling object from your rear view mirror
1
Expert's answer
2020-02-12T10:49:14-0500

a) Calculate the angle the object makes with the vertical direction while speeding up. Is it towards the front or the back of the car?

"\tan{\theta_1}=\frac{a_1}{g}=\frac{3.4}{9.8}"

"\theta_1=19\degree"

It is towards the back of the car.


b) Calculate the horizontal distance from the vertical direction the object is while slowing down. Is this distance away from the vertical towards the front or the back of the car?



"\tan{\theta_2}=\frac{a_2}{g}=\frac{5}{9.8}"

"\theta_2=27\degree"


"x=l\sin{27}=0.3\sin{27}=0.14\ m"

 It is towards the front of the car.